# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ 5.1. Probabilistic and statistical mechanical setup
$\nu^{(N)}=i^{N^{2}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
@@
When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 5.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed). It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
@@ 5.2. Motivation: A conjectural Large Deviation Principle
Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 5.11. Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
@@
showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 5.7 is an equality. As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [42] for the case $\Delta=0$ with $X$ smooth and [77, 79] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
Next, to prove that Conjecture 5.4 A implies Conjecture 1.5, assume that $(X,\Delta)$ admits a KE metric. It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 5.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.12 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric). As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large. This means that $\mu_{-1}^{(N)}(=\mu_{\boldsymbol{m}_{N}}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional. But this is shown in Theorem 3.7.
@@
It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)). If inequality 5.9 can be strengthened as follows, then the second assumption implies the first:
@@
The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds). The second statement is shown in a similar way, using Prop 5.11 below. ∎
@@ 5.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
But this follows from combining the inequalities 5.12 below with the first identity in Proposition 5.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
@@
The main virtue of the following proposition, as compared with Prop 5.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
@@
Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 5.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small. Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 5.10. As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite. It will thus be enough to show that
@@
Hence, the desired formula 5.13 follows from formula 5.14. ∎
@@
Formula 5.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
@@
To this end it is enough to show that the projection $p$ in 5.16 satisfies
@@
Finally, Prop 5.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
@@ 6. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11). Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 5.20, which implies the lemma. ∎
@@
By Prop 5.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 5.4) may be expressed as
@@
Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 6.1, applied to $N=3).$ Thus formula 6.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 5.9. Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant. In fact, taking $e_{1},e_{2}$ and $e_{3}$ as the standard generators of the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ corresponding to the holomorphic vector fields $\partial/\partial z,$ $z\partial/\partial z$ and $z^{2}\partial/\partial z$ on $\mathbb{C}$ a direct computation gives
@@
Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 5.4.1. Thus, accepting formula 6.3, we conclude that
@@
which implies the desired formula 6.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
Finally, as next described, the deferred proof of formula 6.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 6.2). First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
@@
which concludes the proof of formula 6.3.
@@ 6.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 5.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
@@
Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above. Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 6.4. Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@ 6.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 6.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial). By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
@@
(as follows from formula 6.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
@@
Combining this convergence with the convergence 6.5 gives the following explicit formula
@@
which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 6.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 6.6. Note that
@@
where the error terms are positive. Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 6.8:
@@
using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 6.11. The nominator in the second factor in formula 6.8 is handled in essentially the same manner, giving
@@
for $w_{i}=0$ and $V=2.$ The proof of formula 6.6 is thus concluded by invoking formula 6.10.
@@
for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e. $\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 6.13.

### v2

@@ 6.1. Probabilistic and statistical mechanical setup
$\nu^{(N)}=\frac{i^{(Nn)^{2}}}{2^{Nn}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
@@
When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 6.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed). It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
@@ 6.2. Motivation: A conjectural Large Deviation Principle
Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.11. Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
@@
showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.7 is an equality. As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [42] for the case $\Delta=0$ with $X$ smooth and [77, 79] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
Next, to prove that Conjecture 6.4 A implies Conjecture 1.6, assume that $(X,\Delta)$ admits a KE metric. It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.14 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric). As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large. This means that $\mu_{-1}^{(N)}(=\mu_{\boldsymbol{m}_{N}}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional. But this is shown in Theorem 3.7.
@@
It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)). If inequality 6.9 can be strengthened as follows, then the second assumption implies the first:
@@
The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds). The second statement is shown in a similar way, using Prop 6.11 below. ∎
@@ 6.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
But this follows from combining the inequalities 6.12 below with the first identity in Proposition 6.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
@@
The main virtue of the following proposition, as compared with Prop 6.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
@@
Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small. Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.10. As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite. It will thus be enough to show that
@@
Hence, the desired formula 6.13 follows from formula 6.14. ∎
@@
Formula 6.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
@@
To this end it is enough to show that the projection $p$ in 6.16 satisfies
@@
Finally, Prop 6.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
@@ 7. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11). Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.20, which implies the lemma. ∎
@@
By Prop 6.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.4) may be expressed as
@@
Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 7.1, applied to $N=3).$ Thus formula 7.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.9. Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant. In fact, taking $e_{1},e_{2}$ and $e_{3}$ as the standard generators of the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ corresponding to the holomorphic vector fields $\partial/\partial z,$ $z\partial/\partial z$ and $z^{2}\partial/\partial z$ on $\mathbb{C}$ a direct computation gives
@@
Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.4.1. Thus, accepting formula 7.3, we conclude that
@@
which implies the desired formula 7.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
Finally, as next described, the deferred proof of formula 7.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 7.2). First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
@@
which concludes the proof of formula 7.3.
@@ 7.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
@@
Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above. Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 7.4. Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@ 7.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 7.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial). By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
@@
(as follows from formula 7.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
@@
Combining this convergence with the convergence 7.5 gives the following explicit formula
@@
which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 7.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 7.6. Note that
@@
where the error terms are positive. Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 7.8:
@@
using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 7.11. The nominator in the second factor in formula 7.8 is handled in essentially the same manner, giving
@@
for $w_{i}=0$ and $V=2.$ The proof of formula 7.6 is thus concluded by invoking formula 7.10.
@@
for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e. $\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 7.13.